# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ D.3 Model-based Learning Protocol
We use the learning protocol introduced in Algorithm 1 to train Safe-$\text{M}^{3}$-UCRL. For hyperparameters of the model, see Table 1.
@@
To optimize the objective in Section 4 in the subroutine in Line 3, we use MF-BPTT (see Section C.3.1). We parametrize the policy via a fully-connected neural network with two hidden layers of 256 neurons and Leaky-ReLU activations. The output layer returns the agents’ actions using Tanh activation. We use Xavier uniform initialization [21] to randomly initialize weights while we set bias terms to zero. We use 20,000 training epochs with the early stopping if the policy does not improve at least 0.5% within 500 epochs. To prevent gradient explosion, we use L2-norm gradient clipping with max-norm set to 1. Note that in our experiments, a single neural network had enough predictive power to represent the whole policy profile $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})$, but using $T$ networks, one for each individual policy $\pi_{t}$ is a natural extension. For further details about hyperparameters, see Table 2. Additionally, note that we parametrize policy by a neural network with Lipschitz continuous activations (Leakly-ReLU and Tanh). In the case of the bounded neural network’s weights, policy satisfies the Lipschitz continuity assumption (see Assumption 2). In practice, we bound the weights via L2-regularization.
@@
We estimate the confidence set of dynamics functions $\mathcal{F}_{n}$ in the subroutine in Line 5 using deep probabilistic ensemble (see Section C.1). We use an ensemble of 10 fully-connected neural networks with two hidden layers of 16 neurons and Leaky-ReLU activations. The output layer returns the mean and covariance of the confidence set; the mean uses linear activation, and the variance uses SoftPlus activation. We minimize the negative log-likelihood (NLL) for each neural network under the assumption of heteroscedastic Gaussian noise as described in [30]. We randomly split the data from the replay buffer into a training set (90%) and a validation set (10%). We use 10,000 training epochs with the early stopping if the performance on the validation set does not improve for at least 0.5% within 100 consecutive epochs. For further details about hyperparameters, see Table 3.
D.4 Performance Evaluation in the Infinite Regime
@@
In this section, we extend the experiments presented in Section 5. We assume that the number of taxis $m\rightarrow+\infty$ and show two important results. First, we show a conservative behavior of Safe-$\text{M}^{3}$-UCRL by training model with Shannon entropy safety constraint (see Remark 2) with the safety threshold $C$ set to $p=50\%$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$. We then evaluate its performance against a much higher safety threshold, i.e., against the safety threshold induced by $p=85\%$. Figure 4(a) shows that the model never violates stricter safety constraint regardless of its training in a weaker setting. Second, we show the data efficiency of the learning protocol (Algorithm 1) by training models with access to one, five, and ten representative agents for data collection. Figure 4(b) shows that the model with unknown dynamics converges to the model trained with known dynamics almost 6 times faster when using ten representative agents instead of one representative agent. It is a very useful result that can be utilized in many applications. For example, in most transportation applications using tens or even hundreds of representative agents often does not cause cost-related issues. On the contrary, it might be more cost-effective to have ten representative agents for a month than one representative agent for a year.
(a) Safe-$\text{M}^{3}$-UCRL conservative behavior
@@
Figure 5: Safe-$\text{M}^{3}$-UCRL conservative behavior and data efficiency. We train 10 randomly initialized policy profiles where each setup uses the entropic safety constraint $h_{C}(\cdot)\geq 0$. We set the safety threshold $C$ as a percentage $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$ with $k=25$. In (a), the policy profiles trained for satisfying $h_{0.5\log(k^{2})}(\cdot)\geq 0$ never violate $h_{0.85\log(k^{2})}(\cdot)\geq 0$, which shows the conservative behavior of our model. In (b), we show the data efficiency of the learning protocol (Algorithm 1) by comparing learning curves observed when training models that satisfy $h_{0.85\log(k^{2})}(\cdot)\geq 0$ when using one, five and ten representative agents (RA) for data collection. We see that the model trained with 1-RA converges to the performance of the model with known dynamics in around 80 episodes, while it takes 25 and 15 episodes for 5-RA and 10-RA models, respectively. Note that we use log-reward to emphasize learning speeds on a comparable scale.
D.5 Performance Evaluation in the Finite Regime
In this section, we assume a finite number of taxis $m<+\infty$ and approximate the mean-field distribution $\mu_{t}(s)=\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{t}^{(i)}=s)$ with the empirical distribution as explained in Section D.5. The policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to reposition each of $m$ individual taxis in the fleet. In Figure 5(a), we show the relationship between model performance and the number of taxis in the system. As expected, we observe that increasing the number of taxis leads to better performance due to the increased precision of mean-field distribution $\mu_{t}$ approximation at step $t$. By performing 100 randomly initialized runs for each of various fleet sizes, we see that Safe-$\text{M}^{3}$-UCRL learned with unknown dynamics and applied in the finite regime converges to the solution achieved in the infinite regime with known dynamics. Furthermore, the model performs very well already for a fleet of $10^{4}$ taxis, which is in the order of magnitude of the fleet size that operates in Shenzhen. Concretely, in 2016 the fleet had around 17 thousand taxis, with an increasing trend. We also observe that in the finite regime, the difference in performance between Safe-$\text{M}^{3}$-UCRL trained with known and unknown dynamics becomes insignificant. To showcase the practical usefulness of the algorithm, in Figure 5(b), we instantiate 10,000 taxis and display their positions at step $T=12$ after the final repositioning. We observe that the majority of taxis are repositioned to areas of high demand. At the same time, some of them are sent to residential zones in the northwest and northeast to enforce accessibility. It is important to note that some taxis are repositioned to aquatic areas and areas without infrastructure due to two reasons. First, our model guarantees global safety without explicit guarantees for individual/local safety. Notice that undesirable areas might be avoided by safety constraint shaping, e.g., by setting the weight function for these areas to zero as discussed in Section B.2. Second, the model loses some of its accuracy due to the finite regime approximation errors. In Figure 7, we show only a slight decrease in dispersion when Safe-$\text{M}^{3}$-UCRL finite regime approximation is compared to the infinite regime performance.
@@ D.5 Performance Evaluation in the Finite Regime
Taxis’ locations after repositioning action
@@
Figure 6: Safe-$\text{M}^{3}$-UCRL performance in the finite regime. We instantiate a finite number of taxis, each following a policy profile $\bm{\pi}_{n}^{*}$ learned in the infinite regime and satisfying the entropic safety constraint $h_{C}(\cdot)\geq 0$ for $C=0.85\log(k^{2})$ with $k=25$. We perform 100 randomly initialized runs for each of the various fleet sizes. In (a), we see that increasing the number of taxis leads to better performance due to the increased precision of mean-field distribution approximation. Further, we see that Safe-$\text{M}^{3}$-UCRL learned with unknown dynamics and applied in the finite regime converges to the solution achieved in the infinite regime with known dynamics. We also see that the curves for Safe-$\text{M}^{3}$-UCRL trained with unknown and known dynamics almost overlap, i.e., the value of knowing dynamics has little to no insignificance in the finite regime. In (b), we showcase the performance for the realistic number of taxis operating in Shenzhen (10,000 to 20,000). We display the positions of a randomly chosen subset of 1,000 taxis (out of 10,000) at step $T=12$ after the final repositioning. We observe that the majority of taxis are repositioned to areas of high demand, while some of them are sent to residential zones in the northwest and northeast to enforce accessibility.
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
in the infinite regime
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
in the finite regime
Figure 7: Mean-field approximation error in the finite regime. We show the difference in the performance introduced by approximating mean-field distribution with the finite number of taxis as described in Section D.2. In (a), we see that the dispersion of Safe-$\text{M}^{3}$-UCRL in the infinite regime is 96% as elaborated in Figure 2. In (b), the policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to control a fleet of 10,000 taxis. We observe a slight decrease in the dispersion from 96% to 92% due to the mean-field distribution approximation errors.
@@
Table 1: Learning protocol hyperparameters
@@
Table 2: Policy hyperparameters
@@
covariance output activation SoftPlus
@@
batch size 8-128
Table 3: Probabilistic ensemble hyperparameters

### v2

@@ D.3. Model-Based Learning Protocol
We use the learning protocol introduced in Algorithm 1 to train Safe-M3-UCRL. For hyperparameters of the model, see Table 1.
@@
To optimize the objective in Section 4 in the subroutine in Line 3, we use MF-BPTT (see Section C.3.1). We parametrize the policy via a fully-connected neural network with two hidden layers of 256 neurons and Leaky-ReLU activations. The output layer returns the agents’ actions using Tanh activation. We use Xavier uniform initialization Glorot and Bengio (2010) to randomly initialize weights while we set bias terms to zero. We use 20,000 training epochs with the early stopping if the policy does not improve at least 0.5% within 500 epochs. To prevent gradient explosion, we use L2-norm gradient clipping with max-norm set to 1. Note that in our experiments, a single neural network had enough predictive power to represent the whole policy profile $\bm{\pi}{}={}(\pi_{0},\ldots,\pi_{T-1})$, but using $T$ networks, one for each individual policy $\pi_{t}$ is a natural extension. For further details about hyperparameters, see Table 2. Additionally, note that we parametrize policy by a neural network with Lipschitz continuous activations (Leakly-ReLU and Tanh). In the case of the bounded neural network’s weights, policy satisfies the Lipschitz continuity assumption (see Assumption 2). In practice, we bound the weights via L2-regularization.
@@
We estimate the confidence set of transitions $\mathcal{F}_{n-1}$ in the subroutine in Line 5 using a probabilistic neural network ensemble (see Section C.1). We use an ensemble of 10 fully-connected neural networks with two hidden layers of 16 neurons and Leaky-ReLU activations. The output layer returns the mean vector and variance vector (because of the covariance matrix diagonality assumption) of the confidence set; the mean uses linear activation, and the variance uses Softplus activation. We minimize the negative log-likelihood (NLL) for each neural network under the assumption of heteroscedastic Gaussian noise as described in Lakshminarayanan et al. (2017). We randomly split the data from the replay buffer into a training set (90%) and a validation set (10%). We use 10,000 training epochs with the early stopping if the performance on the validation set does not improve for at least 0.5% within 100 consecutive epochs. For further details about hyperparameters, see Table 3.
D.4. Performance Evaluation in the Infinite Regime
@@
In this section, we extend the experiments presented in Section 5.2. We assume that the number of vehicles $m\rightarrow{}+{}\infty$ and show two important results. First, we show a conservative behavior of Safe-M3-UCRL by training model with Shannon entropy safety constraint (see Remark 5) with the safety threshold $C$ set to $p=0.50$ of the maximum Shannon entropy, i.e., $C{}={}p\log(k^{2})$. We then evaluate its performance against a much higher safety threshold, i.e., against the safety threshold induced by $p=0.85$. Figure 7(a) shows that the model never violates stricter safety constraint regardless of its training in a weaker setting. Second, we show the data efficiency of the learning protocol (Algorithm 1) by training models with access to one, five, and ten representative agents for data collection. Figure 7(b) shows that the model under unknown transitions converges to the model trained under known transitions almost 6 times faster when using ten representative agents instead of one representative agent. It is a very useful result that can be utilized in many applications. For example, in most transportation applications, using tens or even hundreds of representative agents often does not cause cost-related issues. On the contrary, it might be more cost-effective to have ten representative agents for a month than one representative agent for a year.
D.5. Performance Evaluation in the Finite Regime
In this section, we assume a finite number of vehicles $m<{}+{}\infty$ and approximate the mean-field distribution
$\mu_{t}(s){}={}\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{t}^{(i)}{}={}s)$
with the empirical distribution as explained in Section D.5. The policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to reposition each of $m$ individual vehicles in the fleet. In Figure 8(a), we show the relationship between model performance and the number of vehicles in the system. As expected, we observe that increasing the number of vehicles leads to better performance due to the increased precision of mean-field distribution $\mu_{t}$ approximation at step $t$. By performing 100 randomly initialized runs for each of various fleet sizes, we see that Safe-M3-UCRL learned under unknown transitions and applied in the finite regime converges to the solution achieved in the infinite regime under known transitions. Furthermore, the model performs very well already for a fleet of 10,000 vehicles, which is in the order of magnitude of the fleet size that operates in Shenzhen. Concretely, in 2016, the fleet had around 17,000 vehicles, with an increasing trend. We also observe that in the finite regime, the difference in performance between Safe-M3-UCRL trained under known and unknown transitions becomes insignificant. To showcase the practical usefulness of the algorithm, in Figure 8(b), we instantiate 10,000 vehicles and display their positions after the final repositioning at step $T=12$. We observe that the majority of vehicles are repositioned to areas of high demand. At the same time, some of them are sent to residential zones in the northwest and northeast to enforce accessibility. It is important to note that some vehicles are repositioned to aquatic areas and areas without infrastructure due to two reasons. First, our model guarantees global safety without explicit guarantees for individual/local safety. Notice that undesirable areas might be avoided by safety constraint shaping, e.g., by setting the weight function for these areas to zero as discussed in Section B.2. Second, the model loses some of its accuracy due to the finite regime approximation errors. In Figure 10, we observe only a slight decrease in dispersion when Safe-M3-UCRL finite regime approximation is compared to the infinite regime performance. To conclude, we showcase the potential of Safe-M3-UCRL for vehicle repositioning in the finite regime, which might be a positive signal for real-world practitioners.
Appendix E Experiments – Swarm Motion
In this section, we extend the swarm motion experiments discussed in Section 5.1 and complement the vehicle repositioning experiments elaborated in Section 5.2 and Appendix D. For this experiment, we use the same private cluster and approximately the same amount of computational resources as reported in Appendix D.
E.1. Modeling Assumptions and Model Parameters
We model the state space $\mathcal{S}$ as the unit torus on the interval $[0,1]$ and set the action space as the interval $\mathcal{A}{}={}[-7,7]$ due to the knowledge of the range of actions from the continuous-time analytical solution Almulla et al. (2017). We approximate the continuous-time swarm motion by partitioning unit time into $T=100$ equal steps of length $\Delta t{}={}1{}/T$. The next state $s_{n,t{}+{}1}{}={}f(z_{n,t}){}+{}\varepsilon_{n,t}$ is induced by the unknown transitions $f(z_{n,t}){}={}s_{n,t}{}+{}a_{n,t}\Delta t$ with $\varepsilon_{n,t}\sim\text{N}(0,\Delta t)$ for all episodes $n$ and steps $t$. The reward function is defined by $r(z_{n,t}){}={}\phi(s_{n,t})-\frac{1}{2}a_{n,t}^{2}-\log(\mu_{n,t})$, where the first term $\phi(s){}={}2\pi^{2}(\sin(2\pi s)-\cos^{2}(2\pi s)){}+{}2\sin(2\pi s)$ determines the positional reward received at the state $s$ (see Figure 5), the second term defines the kinetic energy penalizing large actions, and the last term penalizes overcrowding. Note that the optimal solution for continuous time setting, $\Delta t\rightarrow 0$, can be obtained analytically. Namely, for the infinite time horizon, i.e., $T\rightarrow\infty$, we have
[…]
where $\pi^{*}$ and $\mu^{*}$ form an ergodic solution satisfying $\mu^{*}{}={}U(\mu^{*},\pi^{*},f)$. We use $\mu^{*}$ as a benchmark but note that it might no longer be an optimal solution in the discrete-time setting. Therefore, discrete-time solutions obtained under known transitions serve as a good benchmark for Safe-M3-UCRL performance under unknown transitions. To control overcrowding, we use the Shannon-entropic constraint introduced in Equation 13 instead of having the overcrowding penalty term $\log(\mu_{n,t})$ in the reward. As discussed in Section D.2, Shannon entropy is used because of the discrete mean-field distribution representation. Since higher entropy translates into less overcrowding, we can upfront determine and upper-bound the acceptable level of overcrowding by setting a desirable threshold $C$. Similarly to the discussion in Section D.2, we represent the mean-field distribution by discretizing the state space into $k=100$ uniform intervals and assigning the probability of the representative agent residing within each of them. To compute the mean-field transitions $U(\cdot)$, we use one-dimensional equivalent of Equation 14 and Equation 15. We set the safety threshold $C$ as a proportion $p\in[0,1]$ of the maximum Shannon entropy Equation 13, i.e., $C{}={}p\log(k)$. We initialize safe mean-field distributions $\mu_{n,0}$ as uniform distributions since they maximize Shannon entropy, which makes them safe for every threshold $C$.
E.2. Model-Based Learning Protocol
We follow the same procedure as described in Section D.3 with the hyperparameters from Tables 4 to 6. The only difference compared to Section D.3 is the increase in the complexity of the computational graph because of the high number of steps $T$. Therefore, we use batch normalization Ioffe and Szegedy (2015) to prevent vanishing gradients.
E.3. Performance Evaluation
In this section, we complement the results shown in Section 5.1. We first visualize the positional reward $\phi(\cdot)$ in Figure 5 for the ease of interpretation of the obtained results. We see that the reward has two local maxima, but due to a significant difference in their value, unconstrained benchmarks ignore the lower maxima. On the other hand, Safe-M3-UCRL for $p=0.95$, and to a certain extent for $p=0.5$, take advantage of it by reducing the kinetic energy in the neighborhood of lower maxima as shown in Figure 1(e).
Figure 5. Swarm motion positional reward $\phi(\cdot)$.
Swarm motion positional reward.
Figure 5. Swarm motion positional reward $\phi(\cdot)$.
In Figure 11, we show the mean-field distributions progression over time guided by policies learned by Safe-M3-UCRL for $p=0.5$ and $p=0.95$. We observe that for $p=0.5$, we reach near-stationary distribution after only 10 steps (see Figure 10(d)), i.e., the distribution remains the same until the algorithm terminates at $T=100$. For $p=0.95$, we reach stationarity even faster, as shown in Figure 10(e). The learning process presented in Figure 1(a) is explained by the reduction of the epistemic uncertainty in the estimated transitions $\tilde{f}_{n-1}$. Before the first episode $n=1$, the statistical model is estimated only from trajectories collected by randomly initialized policy $\pi_{0}$. Due to the high epistemic uncertainty in regions that random policy did not explore, upper-confidence hallucinated transitions Equation 8 do not approximate well true transitions (see Figure 10(a)). By episode $n=5$, the model already has a good approximation of the transitions (see Figure 10(b)), while at episode $n=50$, the transitions are known with near-certainty (see Figure 10(c)). These results coincide with the observation in Figure 1(a) where around episode $n=50$, Safe-M3-UCRL starts obtaining the results as if the transitions were known. Note that we implement the toroidal state space on $\mathcal{S}=[0,1]$ by assuming a sufficiently large extension, e.g., $[-1,2]$, of the interval over its borders such that a new state resulting from any possible action is captured with high probability. The new state is then mapped back to interval $[0,1]$ using the modulo operation. For completeness of the analysis, in Figure 6, we show that Safe-M3-UCRL for $p=0.5$ converges to the result obtained under known transitions.
Figure 6. Safe-M3-UCRL learning curve for $p=0.5$ for swarm motion.
Swarm motion learning curve for p equals 0.5.
Figure 6. Safe-M3-UCRL learning curve for $p=0.5$ for swarm motion.
We further validate the observation presented in Figure 1(b) that the constraint for $p=0.5$ results in similar overcrowding as the reward penalty term $-\log(\mu)$. Namely, in Figure 7, we see that Safe-M3-UCRL for $p=0.5$ and M3-UCRL with overcrowding penalty term satisfy the safety constraint $h_{C}(\mu)=0.5\log(k)$ with a similar margin.
Figure 7. Swarm motion safety for $p=0.5$.
Swarm motion constraint violation for p equals 0.5
Figure 7. Swarm motion safety for $p=0.5$.
Figure 8. We showcase Safe-M3-UCRL conservative behavior and data efficiency by training 10 randomly initialized policy profiles and statistical models where each setup uses the entropic safety constraint $h_{C}(\cdot){}\geq{}0$. We set the safety threshold $C$ as a proportion $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$ with $k=25$. In (a), the policy profiles trained for satisfying $h_{0.5\log(k^{2})}(\cdot){}\geq{}0$ never violate $h_{0.85\log(k^{2})}(\cdot){}\geq{}0$, which shows the conservative behavior of our model. In (b), we show the data efficiency of the learning protocol (Algorithm 1) by comparing learning curves observed during training models that satisfy $h_{0.85\log(k^{2})}(\cdot){}\geq{}0$ when using one, five and ten representative agents (RA) for data collection. We see that the model trained with 1-RA converges to the performance of the model under known transitions in around 80 episodes, while it takes 25 and 15 episodes for 5-RA and 10-RA models, respectively. Note that we use log-reward to emphasize learning speeds on a comparable scale.
(a) Safe-M3-UCRL conservative behavior
@@ E.3. Performance Evaluation
Learning protocol conservative behavior and data efficiency.
Figure 8. We showcase Safe-M3-UCRL conservative behavior and data efficiency by training 10 randomly initialized policy profiles and statistical models where each setup uses the entropic safety constraint $h_{C}(\cdot){}\geq{}0$. We set the safety threshold $C$ as a proportion $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$ with $k=25$. In (a), the policy profiles trained for satisfying $h_{0.5\log(k^{2})}(\cdot){}\geq{}0$ never violate $h_{0.85\log(k^{2})}(\cdot){}\geq{}0$, which shows the conservative behavior of our model. In (b), we show the data efficiency of the learning protocol (Algorithm 1) by comparing learning curves observed during training models that satisfy $h_{0.85\log(k^{2})}(\cdot){}\geq{}0$ when using one, five and ten representative agents (RA) for data collection. We see that the model trained with 1-RA converges to the performance of the model under known transitions in around 80 episodes, while it takes 25 and 15 episodes for 5-RA and 10-RA models, respectively. Note that we use log-reward to emphasize learning speeds on a comparable scale.
Figure 9. To showcase Safe-M3-UCRL performance in the finite regime, we instantiate a finite number of vehicles, each following a policy profile $\bm{\pi}_{n}^{*}$ learned in the infinite regime and satisfying the entropic safety constraint $h_{C}(\cdot){}\geq{}0$ for $C=0.85\log(k^{2})$ with $k=25$. We perform 100 randomly initialized runs for each of the various fleet sizes. In (a), we see that increasing the number of vehicles leads to better performance due to the increased precision of mean-field distribution approximation. Further, we see that Safe-M3-UCRL learned under unknown transitions and applied in the finite regime converges to the solution achieved in the infinite regime under known transitions. We also see that the curves for Safe-M3-UCRL trained under unknown and known transitions almost overlap, i.e., the value of knowing transitions has little to no insignificance in the finite regime. In (b), we showcase the performance for the realistic number of vehicles operating in Shenzhen (10,000 to 20,000). We display the positions of a randomly chosen subset of 1,000 vehicles (out of 10,000) at step $T=12$ after the final repositioning. We observe that the majority of vehicles are repositioned to areas of high demand, while some of them are sent to residential zones in the northwest and northeast to enforce accessibility.
@@
Vehicles’ locations after repositioning action
@@
The proposed algorithm performance in finite regime.
Figure 9. To showcase Safe-M3-UCRL performance in the finite regime, we instantiate a finite number of vehicles, each following a policy profile $\bm{\pi}_{n}^{*}$ learned in the infinite regime and satisfying the entropic safety constraint $h_{C}(\cdot){}\geq{}0$ for $C=0.85\log(k^{2})$ with $k=25$. We perform 100 randomly initialized runs for each of the various fleet sizes. In (a), we see that increasing the number of vehicles leads to better performance due to the increased precision of mean-field distribution approximation. Further, we see that Safe-M3-UCRL learned under unknown transitions and applied in the finite regime converges to the solution achieved in the infinite regime under known transitions. We also see that the curves for Safe-M3-UCRL trained under unknown and known transitions almost overlap, i.e., the value of knowing transitions has little to no insignificance in the finite regime. In (b), we showcase the performance for the realistic number of vehicles operating in Shenzhen (10,000 to 20,000). We display the positions of a randomly chosen subset of 1,000 vehicles (out of 10,000) at step $T=12$ after the final repositioning. We observe that the majority of vehicles are repositioned to areas of high demand, while some of them are sent to residential zones in the northwest and northeast to enforce accessibility.
Figure 10. We show the difference in the performance introduced by approximating mean-field distribution with the finite number of vehicles as described in Section D.2. In (a), we see that the dispersion of Safe-M3-UCRL in the infinite regime is $p=0.96$ as elaborated in Figure 3. In (b), the policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to control a fleet of 10,000 vehicles. We observe a slight decrease in the dispersion from $p=0.96$ to $p=0.92$ due to the mean-field distribution approximation errors.
Safe-M3-UCRL under unknown transitions in the infinite regime
Safe-M3-UCRL under unknown transitions in the finite regime
Mean-field approximation error in the finite regime.
Figure 10. We show the difference in the performance introduced by approximating mean-field distribution with the finite number of vehicles as described in Section D.2. In (a), we see that the dispersion of Safe-M3-UCRL in the infinite regime is $p=0.96$ as elaborated in Figure 3. In (b), the policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to control a fleet of 10,000 vehicles. We observe a slight decrease in the dispersion from $p=0.96$ to $p=0.92$ due to the mean-field distribution approximation errors.
Figure 11. Mean-field distributions progression over time.
Mean-field distributions progression over time.
Figure 11. Mean-field distributions progression over time.
Table 1. Learning protocol hyperparameters for vehicle repositioning
@@
Table 2. Policy hyperparameters for vehicle repositioning
@@
Table 3. Probabilistic neural network ensemble hyperparameters for vehicle repositioning
@@
variance output activation Softplus
@@
batch size 8-128 Increasing with the replay buffer size
Table 4. Learning protocol hyperparameters for swarm motion
Hyperparameter Value Description
$N$ 200 Number of episodes
$T$ 100 Number of steps
$k$ 100 Number of discretization segments per axis
$\sigma$ 1 Standard deviation of the system noise
$L_{h}$ $10^{-4}$ Lipschitz constant in Section 4
$\lambda$ 15 Log-barrier hyperparameter in Section 4
# of representative agents 1
Table 5. Policy hyperparameters for swarm motion
Hyperparameter Value Description
# of hidden layers 2
# of neurons 16 Number of neurons per hidden layer
hidden activations Leaky-ReLU
output activation Tanh
$\alpha$ $5\cdot 10^{-3}$ Learning rate
$w$ $5\cdot 10^{-4}$ Weight decay
initialization Xavier uniform
bias initialization 0
$n$ 50,000 Number of epochs
early stopping 0.5% If not improved after 100 epochs
Table 6. Probabilistic neural network ensemble hyperparameters for swarm motion
Hyperparameter Value Description
# of ensemble members 10
# of hidden layers 2
# of neurons 16 Number of neurons per hidden layer
hidden activations Leaky-ReLU
mean output activation Linear
variance output activation Softplus
$\alpha$ $5\cdot 10^{-3}$ Learning rate
$w$ $5\cdot 10^{-4}$ Weight decay
$\beta$ 1 Assumption 4 hyperparameter
initialization Xavier uniform
bias initialization 0
$n$ 10,000 Number of epochs
early stopping 0.5% If not improved for 30 consecutive epochs
replay buffer size 10,000
batch size 8-512 Increasing with the replay buffer size